### v2

Abstract
Coordinated online behavior, which spans from beneficial collective actions to harmful manipulation such as disinformation campaigns, has become a key focus in digital ecosystem analysis. Traditional methods often rely on monomodal approaches, focusing on single types of interactions like co-retweets or co-hashtags, or consider multiple modalities independently of each other. However, these approaches may overlook the complex dynamics inherent in multimodal coordination. This study compares different ways of operationalizing the detection of multimodal coordinated behavior. It examines the trade-off between weakly and strongly integrated multimodal models, highlighting the balance between capturing broader coordination patterns and identifying tightly coordinated behavior. By comparing monomodal and multimodal approaches, we assess the unique contributions of different data modalities and explore how varying implementations of multimodality impact detection outcomes. Our findings reveal that not all the modalities provide distinct insights, but that with a multimodal approach we can get a more comprehensive understanding of coordination dynamics. This work enhances the ability to detect and analyze coordinated online behavior, offering new perspectives for safeguarding the integrity of digital platforms.
keywords:
[…]
1 Introduction
The digital age has transformed the way individuals and organizations interact online, giving rise to novel forms of coordination in social and political arenas. Coordinated behavior, a phenomenon where multiple actors engage in synergic actions in pursuit of an intent [1], has become increasingly relevant in the study of online ecosystems. Coordination in online behavior can manifest across a wide spectrum of intentions and outcomes. It may emerge organically or be deliberately orchestrated, and its perceived value often depends on the observer’s perspective. For instance, coordinated efforts are found in activism [2, 3], boycotts [4, 5], and protests [6, 7], which some may view as socially beneficial, while others may see them as disruptive or even harmful. Conversely, coordination can also be employed for goals widely recognized as problematic, such as disinformation campaigns [8, 9, 10], election interference [3, 11], or astroturfing [12]. Even in cases like coordinated botnets [13, 14, 15, 16], which are typically associated with manipulation, there may be applications aligned with legitimate objectives. Ultimately, the evaluation of coordinated behavior – whether deemed harmful or beneficial – is often subjective and context-dependent. Hence, since coordination in social media exhibits a dual nature, the scope behind coordination can vary widely, ranging from benign collective action to malicious intent aimed at manipulation or deception.
[…]
Refer to captionRefer to captionco-favorite
$w_{fav}$
Figure 1: Coordinated online behavior is inherently multimodal, as users can coordinate through different types of actions.
Understanding and detecting coordinated online behavior is critical for maintaining the integrity of digital platforms [17]. Researchers and policymakers alike seek to develop methodologies to identify patterns of coordination, disentangle genuine collective action from inauthentic manipulation, and mitigate the harmful consequences of coordinated inauthentic behavior. Various computational techniques, including network-based analysis [18, 2, 6, 19] and machine learning models [20, 21, 22] have been employed to unveil hidden networks of coordination. Users may engage in a wide range of coordinated actions, which complicates their detection. For example, users might coordinate through synchronized retweets to amplify specific content [11, 23, 24], through likes [25], or by using identical hashtags to dominate trending topics [26, 27]. More covert forms of coordination can include strategic timing of posts [28], the use of similar narratives across multiple accounts, or the deliberate targeting of specific audiences. These varied manifestations highlight the inherent multimodal nature of coordination in social media, where the mode of action plays a crucial role in shaping the dynamics and outcomes of coordinated behavior. Figure 1 shows how coordinated online behavior is inherently multimodal, as users can coordinate through different types of actions. Therefore, it is crucial to design and provide a robust methodology capable of uncovering coordinated behaviors that manifest through diverse user actions.
While network-based methods have proven effective in detecting such behavior, most approaches either focus on a single type of action or treat different actions separately, resulting in the construction of multiple single-layer networks that miss the opportunity to analyze them in an integrated way. Although some studies have begun to explore multimodal interactions, they fail to fully leverage the richness and complexity that such diversity offers [29, 6]. Many of these approaches flatten the layers into a single aggregated network, thereby losing the distinct characteristics of each modality and the interplay between them [30, 31, 26, 2]. The current landscape reveals two key limitations: first, existing multimodal analyses do not truly exploit the potential of integrating different types of actions; second, there is a lack of systematic investigation into how and whether multimodality offers advantages over analyzing each type of action independently. As a result, we lack not only a clear understanding of the value of multimodality in this context, but also effective frameworks or methods to operationalize it properly. Addressing these gaps is essential for advancing our ability to detect and interpret coordinated behavior in a more comprehensive and meaningful way.
For this reason, it is necessary to adopt a comprehensive approach that accounts for the multimodal nature of coordinated behavior. In this work, we propose different ways to operationalize the detection of multimodal coordinated behavior, by incorporating various types of actions and, consequently, different forms of coordination. By integrating multiple modalities – such as retweets, mentions, hashtag usage, and posting patterns – our method provides a holistic perspective on how users coordinate their efforts. This approach not only captures the diversity of coordinated actions but also enables the detection of subtle and complex behaviors that might be overlooked when focusing on a single modality. By bridging these different dimensions of coordination, our method enhances the understanding and identification of coordinated efforts in social media, contributing to more robust and reliable analysis frameworks.
Our framework includes several multimodal detection methods for coordinated behavior and Figure 2 presents examples of possible ways to operationalize multimodality, along with the number of studies in the literature that have adopted each approach. For example, the method frequently used in the literature – flattening the layers into a single network – could be one solution, albeit perhaps not the most effective. Another approach, which considers different coordination networks and actions independently, as is often done, represents the weakest form of multimodality operationalization, as it overlooks the interplay between modalities. On the other end of the spectrum, the strongest approach forces the model to recognize users as coordinated only if they demonstrate coordination across all types of actions. This stricter operationalization fully embraces multimodality but at the cost of increased selectivity. As we can anticipate, there is an inherent tradeoff in leveraging multimodality. While enriching the model by integrating more modalities allows for the detection of complex and nuanced forms of coordinated behavior, increasing the intersection of multimodal conditions too much may overly constrain the analysis, potentially limiting the detection to only a few instances of coordination. Figure 2 illustrates this tradeoff, highlighting the balance between the level of multimodality and constraints loosening in coordinated behavior detection.
lowhighmultimodalityhighlowconstraints loosening monomodal independent layers union flattening multiplex community detection intersection flattening >20 papers >15 papers >4 papers zero papers zero papers
Figure 2: Trade-off between multimodality and constraints loosening in coordinated behavior detection. Increasing the level of multimodality enhances the complexity and richness of the analysis but may also restrict the detection to only the most tightly coordinated actors.
In summary, our approach not only advances the detection of multimodal coordinated behavior but also provides a comprehensive exploration of how to effectively operationalize multimodality, striking a balance between complexity and detectability.
1.1 Research focus
Understanding coordinated online behavior requires careful analysis of interactions across different digital modalities. Existing approaches typically fall into two broad categories: monomodal methods, which analyze coordination based on a single type of interaction (e.g., retweets, hashtags, or URLs sharing), and multimodal methods, which integrate multiple modalities to capture a more comprehensive picture of coordination. However, the comparative effectiveness of these approaches remains an open question. This study explores the contributions of monomodal and multimodal methods to the detection and analysis of coordinated online behavior, with a particular focus on whether different modalities offer complementary insights or largely redundant information. The following research questions guide our investigation:
  • RQ1: What is the unique contribution of each individual data modality to the detection and analysis of coordinated online behavior? This RQ aims to determine whether monomodal approaches – each focused on a single type of interaction, such as synchronized retweets, hashtags, or URLs sharing – produce similar or differing results between one another. By comparing the outcomes of analyses performed separately on each modality, we assess whether these modalities capture redundant or complementary aspects of coordination. For example, if all modalities provide similar information, their individual analyses would converge on consistent findings. However, if different modalities reveal distinct patterns of coordination, this would indicate that each provides unique and potentially critical information for understanding coordinated behavior.
  • RQ2 How do multimodal approaches contribute to the analysis of coordinated online behavior compared to monomodal approaches? Additionally, how do different implementations of multimodality affect the results? Here we explore the added value of multimodal approaches in detecting and analyzing coordinated online behavior compared to monomodal analyses. While monomodal approaches focus on a single type of interaction, multimodal approaches leverage multiple modalities simultaneously, potentially offering a more comprehensive view of coordination patterns. However, multimodality is a flexible concept that encompasses various ways of combining modalities. Different strategies for integrating modalities may highlight different aspects of coordinated behavior and yield varying analytical outcomes. This question thus addresses two key points: the overall benefit of multimodal analyses relative to monomodal ones, and the impact that different implementations of multimodality have on the results.
To address the research questions outlined above, we conduct our analysis on a Twitter dataset. This dataset offers a rich and appropriate context for studying coordinated online behavior due to the platform’s public and interaction-driven nature. It has been previously used in the study of coordination dynamics, although not in a multimodal setting [3, 32, 23, 33].
1.2 Contributions
This work advances the analysis of coordinated online behavior by examining the role of different data modalities, introducing a novel method for detecting multimodal coordination, and comparing monomodal and multimodal approaches. Additionally, we explore different levels of multimodal integration, highlighting key trade-offs in capturing coordination patterns. Below, we outline our main contributions:
  • We propose the first comparative analysis of operationalizations for detecting multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network and capturing interdependencies among different modalities through a community detection algorithm specifically designed for multiplex structures.
  • We analyze the degree of overlap and divergence in the information provided by different data modalities when used individually in monomodal approaches, revealing how each modality contributes unique or redundant insights. This lays the groundwork for understanding how combining multiple modalities enhances the analysis of coordinated online behavior.
  • We compare monomodal and multimodal approaches to evaluate the advantages and limitations of integrating multiple data modalities, providing insights into the overall contribution of multimodal analyses for detecting and understanding coordinated online behavior.
  • We analyze different implementations of multimodality, ranging from weakly integrated approaches with fewer constraints to strongly integrated ones that impose stricter requirements. This exploration sheds light on the trade-off between capturing broader, more inclusive patterns of coordination and identifying only the most tightly coordinated actors. Our findings clarify how varying levels of integration affect the results and offer guidance on balancing these competing priorities in robust multimodal analyses.
2 Related work
Coordination detection methods can be classified into two main categories depending on their underlying approach: network science or machine learning. Network science approaches can be categorized in single-layer or multi-layer ones. The following sections discuss the existing solutions in each category.
2.1 Machine learning-based approaches
This category of methods applies data mining and machine learning techniques to extract pattern and insights from various data types, including text [34, 35, 36], images [37], audio [22], interactions [20, 21, 7, 38], and temporal data [39, 40, 41]. However, only a few of these approaches extract multimodal features from the data or leverage multiple types of user activity. Despite the diversity of machine learning methods – and the absence of a unifying framework comparable to that of network-based approaches – they can generally be grouped into two broad categories: unsupervised methods, which cluster users based on behavioral similarities, and supervised methods, which classify users as coordinated or non-coordinated based on labeled examples.
Among unsupervised approaches, some methods model user activity as temporal point processes. For instance, Sharma et al. [20] introduced AMDN-HAGE, a generative model that combines activity traces (AMDN) with user group estimation (HAGE) to detect synchronized user groups – assuming all detected coordination to be malicious. Similarly, Zhang et al. [21] jointly learn user-group distributions by leveraging embeddings and temporal logic, clustering users through an expectation-maximization algorithm. In a different direction, Erhardt and Albassam [42], Erhardt and Pentland [43] represent user behaviors as interacting Markov chains within a discrete-time stochastic model to capture coordination dynamics.
Supervised methods, on the other hand, rely on labeled data to identify coordinated users. Zhang et al. [38] address cross-platform coordination by using an auxiliary platform with known coordinated users to train neural time series encoders, followed by a multi-layer perceptron to detect similar behavior on a target platform. Mariconti et al. [22] detect coordinated hate attacks on YouTube by extracting features from video metadata, audio, and thumbnails, and feeding them into an ensemble classifier trained on content that had previously been targeted by coordinated raids. A more lightweight approach is proposed by Francois et al. [41], who construct a small ground-truth of organic and inorganic campaigns and classify new ones based on the confidence intervals of network, temporal, and semantic indicators. ì
2.2 Network science-based single-layer approaches
Most of detection methods in literature exploit the so-called coordination network, a graph where nodes represent users and edges represent a co-action, i.e., an action of the same type performed by two users. For instance, two users retweeting the same tweet perform a co-retweet, or posting a post with the same hashtag, a co-hashtag. Generally, the weight of an edge is proportional to the number of co-actions between the two users. Communities of users can be detected in the coordination network to identify groups of users that are more densely connected to each other than to the rest of the network. These communities are interpreted as groups of users that are coordinating their actions [1]. The majority of the works in literature are monomodal approaches, as they focus on a single type of co-action, such as co-retweet [18, 3, 11, 44], co-hashtag [45, 27, 46], co-mention [47], and co-like [25]. Many others construct several coordination network, one for each co-action, to perform independent analysis and comparing the results. For instance, Weber and Neumann [48, 19] analyze co-retweet, co-URL, co-hashtag, co-mention, and co-reply, Yu [49] analyze users posting posts with similar images and videos, and Ng and Carley [30] construct a co-URL and co-tweet coordination network, i.e., respectively, users posting similar posts and posts with the same URLs.
Although, they perform analysis on different types of co-actions, they do not consider the interdependence between them. Indeed, a model based on the construction of a coordination network for each type of co-action does not capture the interplay between different types of co-actions. For instance, Dash and Mitra [50] refers to their analysis as a multimodal analysis, even if they detect the coordinated communities in each layer independently. With respect to Figure 2, this work falls within the independent layers approach.
2.3 Network science-based multi-layer approaches
As already discussed, online coordination is a complex phenomenon that can involve different types of actions, possibly simultaneously. In this context, some works have proposed to model the coordinated behavior as a multiplex coordination network [29, 6]. At a high level, a multiplex network is a type of network that consists of multiple layers, where each layer represents a different type of relationship or interaction between the same set of nodes. In the context of coordinated behavior, each layer represents a different type of co-action, and the multiplex network can capture the interplay between different types of co-actions. However, modeling the coordination as a multiplex network is not sufficient to capture the multimodality of coordinated behavior if the method used to detect coordinated communities does not fully exploit the multiplex nature of the network [31, 51, 26, 2, 30, 52].
Indeed, most of the works declaring to exploit a multiplex network, in fact, flattens the multiplex network into a single-layer network before applying the community detection method – union flattening in Figure 2. This is done by summing the weights of the edges in the different layers [2], by considering the union of the edges in the different layers [30, 31, 26], or by creating a multigraph, where all the layers are merged into a single layer [52]. This flattening step is motivated by the fact that most of the community detection methods are designed to work on single-layer networks, and are not able to handle multiplex networks. However, this flattening step may lead to a loss of information, as it does not take into account the different types of co-actions.
A couple of works leverages a multi-view clustering algorithm, which is able to handle multiplex networks without flattening them [29, 6]. However, the leveraged algorithm, proposed by Cruickshank [53], optimizes the modularity in each layer, not considering the inter-layer modularity. This means that the clusters are not allowed to span between views or layers, and the modularity can be reduced to just the sum of the intra-layer modularities. This approach is different from the method proposed in this work, which allows communities to span multiple layers. Finally, a work by Emeric and Victor [51] proposes a method to detect coordinated behavior in multiplex networks by applying Louvain on each layer and an iterative probabilistic voting consensus algorithm to achieve consensus clustering [51]. Also this work does not allow communities to span multiple layers and firstly optimizes each layer indipendently.
Finally, the main drawback of the aforementioned works is not only the lack of a method that can fully exploit the multimodal nature of a multiplex coordination network, but also the lack of a comprehensive definition and evaluation of what multimodal coordinated behavior is and how its detection can be operationalized. To address this gap, we present the first comparative analysis of different ways to operationalize the detection of multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network. We also investigate the contribution of the multimodality compared to the monomodal approaches. Indeed, it is still an open question whether different modalities of co-actions provide distinct information for the analysis of coordinated behavior, or they yield equivalent results in terms of coordinated community discovered.
3 Methodology
[…]
Figure 3: Different operationalizations of multimodality. The colored shapes are the detected communities, and the dashed boxes show on which network the community discovery algorithm is performed.
This work aims to investigate the role of multimodality in detecting coordinated behavior by systematically comparing monomodal and multimodal approaches. We first assess the contribution of monomodal methods, analyzing how each type of user action independently reveals coordination patterns.We then examine various ways to operationalize multimodality and assess how each contributes beyond what monomodal approaches can achieve. The framework proposed in this work builds on a network science approach, following the standard conceptual model for coordinated behavior detection [1]. This section is structured into three parts: the foundational network science framework behind our methodology, the detection methods for both monomodal and multimodal approaches, and the methodology for conducting the analysis needed to address the research questions.
3.1 Preliminaries
The analytical framework for coordinated behavior detection exploiting network science-based methods can be structured in four steps: user selection, coordination network construction, network filtering, and community detection [1]. The user selection step identifies a subset $U^{\prime}\subseteq U$ of users based on criteria aligned with the analysis goals, often focusing on the most active accounts, since a small fraction of users generates the majority of activity on social platforms [54]. Common strategies include selecting super-producers, who post frequently [55, 29, 28], or super-spreaders, whose content is widely reshared [3, 56, 57]. Then, the second step involves the construction of the coordination network that in the most general case is a multiplex network $G(V,E,W,L)$ [58]. To build $G$, one must define the types of co-actions to consider (e.g., re-shares, mentions, follows, etc.). When multiple co-actions are used, $G$ is a multiplex network with $|L|$ layers, where $L$ is the set of layers, each corresponding to a selected co-action. However, the majority of existing works leverage a single co-action, i.e., the number of layers is $|L|=1$ and $G$ is a single layer network. Each layer $l\in L$ is a monomodal network in $G$ and is an undirected weighted graph $G^{l}(V^{l},E^{l},W^{l})$, where $V^{l}\subseteq U^{\prime}$, $E^{l}$ and $W^{l}$ respectively denote the set of nodes, edges, and weights of layer $l$. Given a layer $l$, an edge $e^{l}_{jk}$ is created if there exists a co-action of type $l$ between two users $(u_{j},u_{k})$. The edge weight $w^{l}_{jk}$ is obtained via a similarity function that computes pairwise user similarities. Network filtering is an optional but important step used to retain only meaningful coordination signals while reducing network size for more efficient analysis. A common strategy removes edges with weights below a threshold and the resulting isolated nodes [26, 30, 28, 51, 59, 44, 47]. Another approach filters based on the timing of co-actions, keeping only those occurring within specific time windows – either adjacent [60, 8, 19, 18] or overlapping [30, 52, 6, 28, 33]. The final step detects coordinated communities through a community discovery algorithm on the coordination network. Multiplex networks are either flattened prior to discovery [61, 48, 19] or analyzed using multiplex-specific algorithms [33, 58, 62].
3.2 Operationalizations of multimodality
Multimodality can be operationalized in various ways. In the following, we formalize the monomodal and several multimodal approaches that will be compared in the experimental section. Figure 3 illustrates these operationalizations, emphasizing the type of network used for community detection, as this step primarily distinguishes the different methods.
Monomodal. This implementation does not operationalize the multimodality, since coordinated behavior is detected by performing the community detection on a single-layer network, representing a single type of co-action. However, it plays a crucial role as a baseline for evaluating and comparing multimodal approaches. The greatest disadvantage of this approach is that it analyzes only one type of co-action, leading to a loss of information. Formally, given a monomodal approach $l\in L$, a community detection algorithm produces a set of communities $C=\{C^{l}_{1},C^{l}_{2},\dots,C^{l}_{k}\}$, where each $C^{l}_{i}\subseteq G^{l}$. In the rest of the work, we refer to this approach as MONO.
Independent layers. This implementation detects coordinated behavior by identifying communities in each layer independently and then performing comparisons between the detected communities. So it leverages multiple monomodal approaches, focusing on each type of co-action independently. This is the most relaxed form of multimodality that completely loses the interplay between the different types of co-actions. For each modality $l\in L$, a community detection algorithm identifies a set of communities, similarly to the monomodal approach. In the rest of the work, we refer to this approach as INDI.
Union flattening. Multimodality is operationalized by flattening the multiplex network into a single-layer network, by considering the union of the edges and nodes of the different layers. The aggregation of the weights can be done in different ways, for instance, the sum of the weights of the edges or the maximum weight of the edges. The choice of the aggregation method depends on the characteristics of the network and the purpose of the analysis. This operationalization implements a logical OR, requiring that edges exist in at least one layer of the network. For this reason, we also refer to it as union flattening, which corresponds to considering users as coordinated if they engage in at least one type of co-action. Formally, the multiplex network is collapsed into a single-layer network $G^{\cup}=(V,E,W^{\cup})$, where $E$ is the union of all intra-layer edges and $W^{\cup}$ is the set of aggregated weights across the layers. A community detection algorithm is then applied on $G^{\cup}$. In the rest of the work, we refer to this approach as UNFL.
Multiplex community detection. Multimodality is operationalized by performing a community discovery algorithm designed for multiplex networks. In the literature, several algorithms have been proposed for this purpose [58], with the advantage of leveraging inter-layer connections, which are implicit in multiplex networks. In such networks, two nodes representing the same user across different layers are implicitly linked, allowing the algorithm to capture multimodal coordination. Formally, a community detection algorithm is applied to the full multiplex network, yielding a set of communities that might span multiple layers. In the rest of the work, we refer to this approach as MULTI.
Intersection flattening. Multimodality is operationalized by flattening the multiplex network into a single-layer network, by considering the intersection of the edges and nodes of the different layers. As well as the union flattening, the aggregation of the weights can be done in different ways. This operationalization represents the strictest form of multimodality and implements a logical AND, requiring that edges appear in all layers of the network. Accordingly, we refer to it as intersection flattening, as it considers users coordinated only if they engage in all types of co-action. Formally, the multiplex network is collapsed into a single-layer network $G^{\cap}=(V,E^{\cap})$, where $E^{\cap}$ is the intersection of all intra-layer edges. A community detection algorithm is then applied on $G^{\cap}$. In the rest of the work, we refer to this approach as INTFL.
3.3 Coordinated communities comparison
In this section, we present the methodology used to address both research questions. This general approach applies to the analysis of the independent layers operationalization – by assessing the contribution of each monomodal approach – and to the comparison between multimodal and monomodal approaches. To compare two operationalizations, our method measures the overlap between the communities they produce. Based on the degree of overlap, we identify whether communities and nodes are detected by only one of the two approaches or by both. This allows us to label communities and nodes as gained, lost, or common.
3.3.1 Overlapping communities
To quantify the degree of overlapping between the communities identified by two different approaches $A$ and $B$, we define the overlap matrix $O$. Given the sets of communities detected by the two approaches $C^{A}=\{C_{1}^{A},\ldots,C^{A}_{k}\}$ and $C^{B}=\{C_{1}^{B},\ldots,C^{B}_{k^{\prime}}\}$, we define the overlap matrix $O=[o_{ij}]\in\mathbb{R}^{k^{\prime}\times k}$ where $o_{ij}$ denotes the harmonic mean overlap at row $i$ and column $j$. We can observe, that the size of the two sets $C^{A}$ and $C^{B}$ might not be the same, i.e, $k^{\prime}\leq k$. Hence, $\forall i=1,\ldots,k^{\prime},j=1,\ldots,k$ we define $o_{ij}$ as:
$o_{ij}^{AB}=\frac{2\cdot r_{ij}^{AB}\cdot r_{ji}^{AB}}{r_{ij}^{AB}+r_{ji}^{AB}}$ (1)
where $r$ is defined as the relative overlapping of a community $i$ with respect to a community $j$. We remark that $r_{ij}^{AB}$ is different from $r_{ji}^{AB}$ and is defined as:
$r_{ij}^{AB}=\frac{|int_{ij}^{AB}|}{|C_{i}^{A}|}$ (2)
where $int_{ij}^{AB}=C_{i}^{A}\cap C_{j}^{B}$ is the intersection between two communities.
In case we need to compare MONO against MULTI, we account for the fact that multiplex communities may span multiple layers. To this end, for each multiplex community $C$ we select only the subgraph $C|_{l}$ involving the subset of nodes belonging to the specific layer $l$, which corresponds to the MONO approach, i.e., we select $C|_{l}=C_{i}\cap V^{l}$.
3.3.2 Lost, common, and gained communities
In this section, we analyze the overlap matrices computed in the previous step to classify communities as lost, common, or gained. To address the first research question, we examine whether the results of the INDI approaches differ or largely overlap, assessing the uniqueness of each modality. Similarly, we compare UNFL and MULTI against INDI. The classification of communities as lost, gained, or common is inherently perspective-dependent, as it relies on which approach is chosen as the reference. When comparing two community detection methods – such as a MONO and a multimodal one – the same set of unmatched communities may be considered either lost or gained depending on the direction of the comparison. In our analysis, we adopt the perspective of the multimodal approach as the reference, since our goal is to evaluate its contribution relative to the monomodal baselines. By doing so, we can properly interpret which communities are newly identified by the multimodal view, which ones are retained from the monomodal ones, and which are no longer detected. This choice of perspective ensures that the labeling captures the added value of multimodal integration. Hence, in the following analysis, we designate the multimodal approaches as the reference method, denoted by approach $B$. Instead, when comparing the INDI approaches, the choice of the reference method is less critical, although a MONO approach must still be selected as the baseline for consistency. For instance, in this work, we adopt the co-retweet modality as the reference.
Formally, the first step is the execution of the Hungarian algorithm [63], which provides the set of the couple of communities maximizing the sum of the harmonic mean overlap. Formally, let $\mathcal{P}=\{1,\dots,k\}\times\{1,\dots,k^{\prime}\}$ be the Cartesian product of the communities of the approach $A$ and $B$, the Hunagrian algorithm finds an assignment $\pi^{*}$ such that the total overlap is maximized, i.e., $\pi^{*}=\arg\max_{\pi\in\mathcal{P}}\sum_{(i,j)\in\pi}o_{ij}$, where $\pi^{*}\subseteq\mathcal{P}$ is a one-to-one matching. We must note that $\pi^{*}$ is a subset, since if $k>k^{\prime}$, not all communities of $C^{B}$ are matched. Each pair $(i,j)\in\pi^{*}$ represents a match between community $A_{i}$ and community $B_{j}$. Hence, we define the non-matching communities of approach $A$ as $A_{\text{unmatched}}=\left\{i\in\{1,\dots,k\}\;\middle|\;\nexists j\in\{1,\dots,k^{\prime}\}\text{ such that }(i,j)\in\pi^{*}\right\}$ and similarly, for approach $B$, $B_{\text{unmatched}}=\left\{j\in\{1,\dots,k^{\prime}\}\;\middle|\;\nexists i\in\{1,\dots,k\}\text{ such that }(i,j)\in\pi^{*}\right\}$.
Given a threshold $\theta\in[0,1]$, we define two functions $\lambda^{lost}$ and $\lambda^{gain}$ to label respectively the sets of communities of $C^{A}$ and $C^{B}$ as follows:
$\lambda^{lost}(C_{i}^{A},O)=\begin{cases}\text{common}&\text{if }\exists j\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}\geq\theta\\ \text{lost}&\text{if }\exists j\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}<\theta\\ \text{lost}&\text{if }i\in A_{\text{unmatched}}\end{cases}$ (3)
$\lambda^{gain}(C_{j}^{B},O)=\begin{cases}\text{common}&\text{if }\exists i\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}\geq\theta\\ \text{gained}&\text{if }\exists i\text{ s.t. }(i,j)\in\pi^{*}\text{ and }o_{ij}^{AB}<\theta\\ \text{gained}&\text{if }j\in B_{\text{unmatched}}\end{cases}$ (4)
3.3.3 Lost, common, and gained nodes
The community labeling methodology described in Section 3.3.2 effectively summarizes the overlap information, highlights differences between approaches, and offers a more intuitive interpretation of the results. However, it relies on an arbitrary threshold $\theta$ to distinguish between lost, common, and gained communities, which may influence the outcomes. To address this limitation, we also propose a node-level labeling method that classifies nodes as lost, common, or gained without requiring any threshold. While this approach avoids the arbitrariness of $\theta$, it only provides insights at the node level and does not capture information about entire communities.
Let $\pi^{*}$ be the optimal assignment obtained via the Hungarian algorithm, $A_{\text{unmatched}}$ the non-matching communities of approach $A$, and $B_{\text{unmatched}}$ for approach $B$. We define a function $\lambda^{node}$ to label the set of nodes $V$ as follows:
$\lambda^{node}(V,C_{i}^{A},C_{j}^{B},O)=\begin{cases}\text{common}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{i}^{A}\cap C_{j}^{B}\\ \text{lost}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{i}^{A}\setminus C_{j}^{B}\\ \text{gained}&\text{if }\exists(i,j)\in\pi^{*}\text{ such that }v\in C_{j}^{B}\setminus C_{i}^{A}\\ \text{lost}&\text{if }\exists i\in A_{\text{unmatched}}\text{ such that }v\in C_{i}^{A}\\ \text{gained}&\text{if }\exists j\in B_{\text{unmatched}}\text{ such that }v\in C_{j}^{B}\end{cases}$ (5)
This classification applies to the same three comparison scenarios as defined for communities.
4 Experimental setting
In this section, we present the dataset utilized for the experiments, detail the preprocessing steps applied, and describe the implementation choices adopted for the method. The proposed method can be applied to any dataset that contains multiple types of social media actions, such as sharing, replying, or publishing posts with hashtags. In this study, we use the dataset from Nizzoli et al. [3], collected via Twitter Streaming APIs during the 2019 UK General Election. The collection spans from 12 November to 12 December 2019, covering the month leading up to the election. It includes all tweets mentioning at least one of the most popular election-related hashtags, as well as all tweets from the official accounts of the two main parties and their leaders, along with the interactions (retweets and replies) those tweets received.
4.1 Preprocessing
Some initial decisions are needed to be taken to define the input dataset for the proposed method. We filter out the hashtags and mentions adopted for the data collection, as they would alter the results of the analysis, since we are searching for common actions. So the following hashtags have been filtered: GE2019, GeneralElection19, GeneralElection19, GeneralElection2019, VoteLabour, VoteLabour2019, ForTheMany, ForTheManyNotTheFew, ChangeIsComing, RealChange, VoteConservative, VoteConservative2019, BackBoris, GetBrexitDone. The following mentions have been filtered: jeremycorbyn, UKLabour, BorisJohnson, Conservatives.
In the literature, the domain is often considered instead of the full URL, as it provides a more generalized and meaningful representation of the source [64, 65, 59]. Following this approach, we extracted the domain from the URLs to focus on this higher-level identifier. Subsequently, the most frequently shared URLs are manually reviewed to filter out [^1] The complete list of domain URLs removed is: a) twitter.com, b) cards.twitter.com, c) youtube.com, d) instagram.com, e) open.spotify.com, f) google.com, g) reddit.com, h) play.google.com, i) bing.com, j) google.co.uk. those that, based on qualitative assessment within the specific context of the dataset, are not considered plausible targets of coordination, aligning with established methodologies and enhancing the interpretability of the analysis [60, 65, 26].
4.2 Implementation choices
In the following, we show the implementation choices regarding the framework presented in Section 3.
Network-science method. In the user selection step, we aim to maximize the number of users while ensuring sufficient overlap across different actions, without incurring high computational costs. To this end, for each action type $l$, we select the top 5% most active users, denoted as $U^{\prime}_{l}$, and define the final user set $U^{\prime}$ as the union of these subsets. While this does not guarantee that each user is active in all modalities, it enables us to retain more users compared to prior studies [3, 56, 57], while ensuring substantial inter-layer overlap and a shared user base for multimodal comparisons.
The second step is the coordination network construction. We focus on the most common actions (co-actions) on Twitter/X, i.e., co-retweet (RTW), co-reply (RPL), co-mention (MEN), co-hashtag (HST), co-URL (URL). We build a multiplex network with five layers, each corresponding to a different co-action. Every user is represented by a vector capturing their activity: retweeted post IDs, replied-to tweet IDs, mentioned user IDs, hashtags, and shared URLs. In each layer, we calculate the edge weight between pairs of users using the cosine similarity of their TF-IDF-weighted vectors [1]. This approach reduces the influence of widely shared or viral content and emphasizes the significance of less common items [3, 57]. We also compute, the absolute overlapping between the vectors representing the users to have a measure of the number of common actions generating the edge. This measure is useful to understand the significance of the edge.
Moreover, in this work, we consider an overlapping time window, evenly distributed in time, of 6 hours with a shift of 5 hour and consequently a 1-hour overlap. This choice is is motivated by the fact that we want to capture both spontaneous and malicious coordination, without focusing on a specific type of behavior. In this way, according to the period of reference (one month), we obtain around 147 time windows (the precise number can vary according to the action). In this work, we build a coordination network for each time window and then we merge them into a single network by averaging the weights of edges across time windows and summing those that appeared in multiple windows. The final coordination network is a multiplex network with 5 layers, one for each type of action.
Network filtering step aims to filter out insignificant edges to highlight meaningful coordination, while preserving as much user overlap across layers as possible. First, edges generated by few common actions are removed, using action-specific thresholds chosen to retain up to $20,000$ nodes per layer, balancing coverage and computational cost. Then, a second filter is applied to edge weights, using the median weight in each layer as a threshold to retain a representative set of edges. This two-step process ensures both relevance and scalability, as detailed in Table 1.
Table 1: Number of nodes and edges in the coordination network before and after the filtering steps. The thresholds are set on the number of common actions generating the edge ($th_{a}$) and on the edge weights ($th_{w}$).
no filter filtered on actions filtered on weight
layer nNodes nEdges $\boldsymbol{th_{a}}$ nNodes nEdges $\boldsymbol{th_{w}}$ nNodes nEdges
RTW $58,473$ $379,668,222$ 26 $19,311$ $3,509,326$ 0.14 $17,200$ $1,426,822$
RPL $33,533$ $38,174,756$ 4 $17,061$ $2,663,906$ 0.43 $16,186$ $1,453,235$
URL $43,300$ $117,456,182$ 3 $19,695$ $4,601,506$ 0.88 $18,159$ $1,494,889$
MEN $61,177$ $1,146,693,369$ 48 $19,929$ $6,120,082$ 0.16 $15,904$ $1,298,155$
HST $58,666$ $382,965,201$ 14 $19,973$ $3,042,391$ 0.25 $18,225$ $1,478,510$
Multimodal operationalizations. The implementations choices for the multimodal operationalizations described in Section 3.2 are the following:
  • INDI is implemented by performing Louvain algorithm [66] on each layer independently. The resolution parameter is set to 1, allowing the detection of communities of varying sizes.
  • UNFL is implemented by flattening the multiplex network into a single-layer network, with 3 different strategies of weight aggregation: 1. UNFL(nw): not weighted edges [31, 26, 30], 2. UNFL(ec): weight equal to the number of layers in which the edge is present, 3. UNFL(sum): weight equal to the sum of the weights of the edges in the different layers [2]. We then apply the Louvain algorithm on the flattened network, setting the resolution parameter to 1.
  • MULTI is implemented by applying the Generalized Louvain algorithm [67] on the multiplex network. This method adapts the Louvain algorithm for multiplex networks and it maximizes modularity across all layers, accounting for both intra-layer and inter-layer structures. A key feature is the coupling parameter $\omega$, which regulates the weight of inter-layer connections. This approach enables communities to span layers without requiring presence in all of them, offering a balanced form of multimodality. In this study, $\omega$ is fixed to enable fair comparison across different multimodal operationalizations. We set the resolution parameter $\gamma=1$ and the coupling parameter $\omega=0.1$.
  • INTFL is implemented by flattening the multiplex network into a single-layer network, considering the intersection of the edges and nodes of the different layers and the weight equal to the sum of the weights of the edges in the different layers. We apply the Louvain algorithm on the flattened network, setting the resolution parameter to 1.
5 Results
In this section, we present the results of the experiments conducted to evaluate the proposed approach. In the first section, we compare the layers of the coordination network, while in the second part we analyze the different ways of operationalizing the multimodal coordinated behavior.
5.1 RQ1: Contribution of modalities
The first research question seeks to determine whether different modalities contribute equally to the understanding of coordinated behavior. If distinct modalities highlight different users or communities, it is evident that they convey unique information. However, even when the same users or communities emerge across modalities, their properties and characteristics may vary significantly. This variation could reveal unique patterns, roles, or dynamics within specific layers, underscoring the importance of analyzing the interplay between modalities to fully capture the multifaceted nature of coordination.
5.1.1 Layers comparison
Table 2: Statistics of the coordination network after applying the filtering steps.
weight statistics
approach mean median stdDev max min nNodes nEdges nCC nCom
monomodal RTW 0.189 0.171 0.060 1 0.126 $17,891$ $1,760,094$ 64 100
RPL 0.610 0.584 0.121 1.000 0.451 $16,071$ $1,334,039$ 6 23
URL 0.923 0.936 0.074 1.000 0.791 $18,741$ $2,302,080$ 117 134
MEN 0.173 0.151 0.072 1.000 0.111 $18,444$ $3,047,793$ 61 95
HST 0.362 0.335 0.101 1.000 0.246 $18,300$ $1,520,067$ 126 186
multimodal UNFL (nw) - - - - - $39,646$ $7,303,637$ 106 175
UNFL (ec) 1.364 1.000 0.688 5.000 1.000 $39,646$ $7,303,637$ 106 154
UNFL (sum) 0.596 0.537 0.371 3.971 0.111 $39,646$ $7,303,637$ 106 169
MULTI - - - - - - - - 180
INTFL 2.245 2.219 0.235 3.849 1.826 257 375 19 27
$\dagger$ nCC: number of connected components, $\ddagger$ nCom: number of communities
Statistics comparison. In Table 2, we report key statistics for each implemented method, including edge weight statistics, the number of nodes, edges, connected components, and communities. UNFL (nw) contains empty values for the weight statistics because the flattened network is unweighted. Similarly, the MULTI approach also presents mostly empty values across the statistics, as it operates on a multiplex network composed of five separate layers – meaning that the relevant statistics are computed individually for each layer. Among the layers, URL and RPL exhibit the highest mean edge weights, while MEN has the lowest. RTW has the highest number of edges, whereas MEN has the fewest. Interestingly, RPL has only six connected components, with one large connected component and a few small isolated ones. However, for all layers, a large connected component exists, which is the primary focus of our analysis. In terms of community count, the layers show similar results, except for HST, which has the highest number of communities, and RPL, which has the lowest.
Therefore, the upper part of Table 2 compares different single-layer approaches, while the lower part presents the results of the multimodal approaches. This analysis offers an initial comparison of the different operationalizations of multimodality. Among them, INTFL results in the lowest number of nodes and edges, whereas UNFL captures the highest. INTFL significantly reduces the network size since it retains only edges present in all layers, identifying the fewest communities. In addition, the mean edge weight is significantly higher with respect to the other flattening approaches. On one hand, this highlights that INTFL captures extremely significant coordinated behaviors, but on the other hand, it is very selective and it is the strictest operationalization of multimodality, limiting the amount of captured information compared to other methods. For this reason, our focus will shift to comparing INDI, UNFL, and MULTI approaches.
(a) Actor coverage.
(b) Edge coverage.
(c) Pearson degree correlation.
Figure 4: Metrics comparison between the layers of the coordination network. The computed metrics are on the actors and edges coverage, Jaccard similarity, Jeffrey divergence, and Pearson correlation.
Metrics comparison. In this analysis, we leverage metrics recommended by Bródka et al. [68] for evaluating the layers of a multiplex network: actor coverage, edge coverage, and Pearson correlation. Figure 4 presents heatmaps comparing the layers, or modalities, of the coordination network. Each heatmap displays a pairwise comparison between two layers, with rows and columns representing the layers and cells containing the computed metrics.
Actor coverage. The coverage metrics assess the percentage of actors and edges shared between two layers, offering insights into the degree of overlap. These metrics are directional, meaning that the coverage $cov(i,j)\neq cov(j,i)$. Specifically, the value in cell $(i,j)$ represents the percentage of actors or edges in layer $i$ that are also present in layer $j$. The coverage values range from 00 to $1$, where $1$ indicates complete overlap between the actors or edges in layer $i$ and those in layer $j$.
Results, showed in Figure 4(a), indicates that MEN has the largest values with both RTW and HST in both directions. This implies that the actors involved in MEN are largely present in RTW and HST, and vice versa. Conversely, RPL exhibits the smallest actor coverage with all other layers, signifying that actors in RPL are less involved in the other layers. URL, excluding RPL, shows a moderate level of actor coverage with all other layers. These findings suggest that RPL is the most distinct layer in terms of actor involvement, while RTW and MEN appear to be quite similar. In particular, the actor coverage values are comparable in both directions, reflecting the similar number of actors present in these two layers.
Edge coverage. Results, showed in Figure 4(b), mirror those of actor coverage but exhibit some directional variations. For example, RTW has an edge coverage of $0.707$ with MEN, while MEN has only $0.409$ with RTW. This disparity arises because co-mention contains twice as many edges as RTW, leading to greater representation of RTW edges within MEN. HST also shares a significant proportion of edges with MEN. In contrast, RPL and URL show the smallest edge coverage with all other layers. For RPL, this is expected due to its low actor coverage. However, the small edge coverage of URL is surprising, given its moderate actor coverage. This discrepancy might result from the unique nature of URL-sharing actions compared to other actions.
Pearson degree correlation. The Pearson correlation evaluates the linear relationship between the degree of actors in two layers, with values ranging from $-1$ to $1$. A correlation of $-1$ indicates that high-degree actors in one layer are low-degree in the other and vice versa, while a correlation of $1$ indicates that high-degree (or low-degree) actors in one layer are also high-degree (or low-degree) actors in the other. Notably, this metric focuses on the number of incident edges for each actor-layer pair, without considering the specific adjacent actors.
Results, showed in Figure 4(c), reveal a strong similarity in actor degree distribution among RTW, MEN, and HST, with the strongest correlation observed between RTW and MEN. In contrast, the weakest correlation is between RPL and URL, further highlighting their distinctiveness.
These preliminary findings highlight that RPL, and to a lesser extent URL, stand out as distinct layers, while RTW, MEN, and HST demonstrate considerable similarity. This result underscores the complementary nature of the different layers: although they provide diverse information, actors tend to maintain similar roles across layers.
5.1.2 Communities comparison
Normalized Mutual Information. We begin with a quantitative analysis to evaluate the similarity between the communities identified by different monomodal approaches. To achieve this, we use Normalized Mutual Information ($NMI$), a metric designed to compare two partitions of a set of nodes [69, 70, 71]. $NMI$ quantifies the extent to which the community structures from different modalities align, providing a standardized measure of agreement between the partitions. This analysis allows us to assess how closely the community detection results from different monomodal approaches correspond to other ones. The $NMI$ is computed considering the communities with more than 200 nodes for all couples of co-actions, and the results are reported in Figure 5. This analysis confirms that RTW and MEN are quite similar in terms of communities detected, reaching the maximum value equal to $NMI=0.38$. The smallest values are reached by all co-actions with RPL, which is in line with the previous results, showing that RPL is the most different co-action in terms of detected communities.
Figure 5: Normalized Mutual Information ($NMI$) heatmap between the set of communities with size greater than 200 for each couple of co-actions.
(a) RTW vs. INDI Lost, common, and gained communities.
(b) RTW vs. INDI Lost, common, and gained nodes.
Figure 6: Bar charts showing the number of the lost, common, and gained communities and nodes switching from INDI to RTW.
Lost, common, and gained communities and nodes. Based on the methodology described in Section 3.3.1, we compute the overlapping of all possible combinations of communities, with size greater than 200 nodes. The matrices obtained from the analysis can be effectively visualized as heatmaps, providing an intuitive representation of the underlying relationships. Appendix A reports the heatmaps for the performed comparisons. Specifically, to answer the first research question RTW vs. INDI, and the second research question UNFL vs. INDI and MULTI vs. INDI. A comparison of the overlapping matrices of the three UNFL strategies reveals minimal differences among them. Therefore, we will focus on the UNFL (sum) strategy in subsequent analyses, as it retains the most information. Given the large number of results to present, we report only the aggregated outcomes for brevity.
Hence, following the methodology defined in Section 3.3.2, we analyze the overlapping matrices of the communities identified by the different approaches, by labeling the communities as lost, common, and gained switching from one approach to another one. We set the threshold $\theta=0.5$, but we obtain similar results for $\theta=0.7$ and $\theta=0.9$. A higher threshold reduces the number of common communities, but the patterns observed remain still valid. We focus only on RTW vs. INDI, since it is just a matter of perspective, and RTW is the most common and studied action in the literature [3, 18, 26, 72, 73]. Similarly, we analyze the lost, common, and gained nodes, according to the methodology defined in Section 3.3.3. Figure 6 presents bar charts summarizing how the communities and nodes identified in the INDI approaches correspond to those found in the RTW network. Specifically, communities and nodes are classified as lost, common, or gained, based on whether they are no longer detected, consistently identified, or newly discovered in RTW compared to the INDI approaches. These results allows to compare the different MONO approaches, allowing to understand if different co-actions provide different information.
Figure 6(a) highlights that RPL and RTW capture completely different structures: they share no communities in common, with RTW identifying five new communities and losing the two detected by RPL. Similarly, URL and RTW exhibit almost entirely distinct results, as RTW loses seven communities compared to URL and identifies four new ones. In contrast, RTW and MEN fully overlap, detecting the same set of coordinated communities and thus conveying identical information about user coordination. Finally, RTW and HST share four communities, while three communities are detected exclusively by HST. This indicates that RTW and HST partially capture complementary aspects of the coordinated behavior, while still identifying some common structures.
Similarly, Figure 6(b) shows that RTW differs substantially from both RPL and URL in terms of the coordinated nodes it detects. However, when shifting the focus from communities to individual nodes, no MONO approach clearly emerges as more informative. Rather, the information gained by RTW is approximately balanced by what it loses, indicating that each modality captures distinct yet equally important facets of coordination. Consistent with previous findings, RTW remains strongly aligned with MEN, while a similar but weaker alignment is observed with HST.
5.1.3 Common communities characterization
While we have compared the community results across modalities in terms of node overlap, this does not ensure that the position or role of these communities within the network remains consistent across layers. Even if the same users are identified as part of the communities, their structural properties and interactions may vary depending on the modality. To address this, we compute a set of metrics to characterize the communities previously labeled as common across modalities. This analysis allows us to assess whether these communities exhibit similar structural and functional characteristics across different layers, providing deeper insights into the interplay between modalities. To conduct this analysis, we compute a set of metrics for each community identified as common across modalities (i.e., detected by both RTW and another MONO approach). The metrics include size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity [74]. These metrics provide a comprehensive characterization of the communities, capturing their structural and functional properties within the network. This approach allows us to evaluate and compare the communities’ characteristics across different modalities systematically.
Table 3: Cosine similarity between the vectors of the communities in common with RTW communities. The last column reports the average cosine similarity for each layer.
$layer_{1}$ $layer_{2}$
layer community layer community cosine similarity mean
RTW 0 MEN 0 0.993 0.971
RTW 1 MEN 1 0.952
RTW 2 MEN 2 0.969
RTW 3 MEN 3 0.968
RTW 4 MEN 4 0.976
RTW 0 HST 0 0.961 0.923
RTW 1 HST 2 0.836
RTW 2 HST 3 0.982
RTW 4 HST 5 0.915
RTW 1 URL 4 0.537 0.537
For each pair of common labeled communities (as identified in Figure 6(a)), we calculate the cosine similarity between their metrics. Higher cosine similarity values indicate greater similarity between the communities in terms of their structural and functional properties. Additionally, we compute the mean cosine similarity for the common communities within each layer. The results, presented in Table 3, reveal that MEN and HST layers exhibit very high levels of cosine similarity, suggesting strong alignment in the structural and functional characteristics of the detected communities. This aligns with the finding that RTW shares a substantial number of communities with both MEN and HST, reinforcing the idea that these modalities capture closely related coordination patterns. In contrast, the URL layer not only has very few communities in common with RTW, but the only shared community also exhibits significant structural differences. This suggests that URL captures a distinct form of coordination, diverging more from retweet-based interactions compared to mention and hashtag.
A more detailed representation of the common communities similarity is presented by Figure 7, which visualizes as a starplot the normalized scores between 0 and 1 of the aforementioned metrics for three examples of couple of common community. The starplot are color-coded according to the modality to which each community belongs, with a distinct color representing each MONO approach. This visualization allows us to compare the metric scores of matched communities and assess the degree of similarity between them. If two modalities convey similar information, the matched communities will exhibit comparable metric scores, resulting in similar starplot shapes. Conversely, differences in the shapes of the radar plots indicate variations in the information captured by the modalities. This representation effectively highlights the similarities and discrepancies across modes in a clear and intuitive manner. As Figure 6(a) shows, RPL and RTW do not have communities in common, and for this reason there is no starplot for comparison. MEN common communities exhibit similar values across all metrics, whereas HST communities are characterized by a high average weight, and their starplot shape slightly differs from that of RTW. Lastly, URL, which shares only one common community with RTW, presents a completely distinct starplot shape, with notably high values for both average weight and assortativity.
(a) RTW vs. MEN
(b) RTW vs. HST
(c) RTW vs. URL
Figure 7: Example of metrics comparison of three couple of common communities between RTW and MEN, HST, and URL.
Figure 8: Visualization of common communities between RTW and other co-actions in a two-dimensional space using PCA. Each community is represented by a vector of metrics, i.e., size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These high dimensional vectors are projected into a two-dimensional space using PCA. Communities are represented with distinct point shapes based on the co-action and are linked to their corresponding RTW community. Each RTW community may be associated with multiple communities from different co-actions, with all connected communities sharing the same color for clarity.
The starplot visualization allows to have a quick understanding of the differences between the communities, but it is very detailed and does not provide a global view of the differences. Hence, we represent each community as a feature vector of the aforementioned metrics. Then the feature vectors are projected into a two-dimensional space using Principal Component Analysis (PCA). Figure 8 shows the scatterplot in two PCA dimensions, where each community is linked with the one in common with RTW and identified with a different point shape according to the co-action. Each RTW community may be associated with multiple communities from different co-actions, and the cluster of connected communities are depicted with the same color. The PCA visualization highlights how each cluster of common communities – same color – occupies a distinct region of the plot. This spatial separation suggests that common communities of different INDI are structurally similar to each other while remaining distinct from non-common communities. This reinforces the idea that coordination patterns within common communities are preserved across different modalities. However, an exception emerges with the URL layer, which appears more dispersed and does not conform to the same clustering pattern. This aligns with previous findings that URL is the most distinct and unconventional layer, capturing a different form of coordination compared to the others.
5.2 RQ2: Multimodal approaches comparison
The second research question seeks to examine how multimodal approaches contribute to the analysis of coordinated online behavior compared to monomodal approaches and how different implementations of multimodality affect the results. Similarly to the first research question, we exploit the definition of lost, common, and gained communities and nodes.
5.2.1 Communities comparison
(a) UNFL (sum) vs. INDI.
(b) MULTI vs. INDI.
Figure 9: Bar charts showing the number of lost, common, and gained communities switching from INDI to UNFL (sum) and MULTI.
(a) UNFL (sum) vs. INDI.
(b) MULTI vs. INDI.
Figure 10: Bar charts showing the number of lost, common, and gained nodes switching from INDI to UNFL (sum) and MULTI.
We categorize communities as lost, common, or gained when exploiting UNFL and MULTI with respect to INDI. Figure 9 presents bar charts illustrating the number of communities across these categories, and Figure 10 the labelling for the nodes.
Multimodal approaches comparison. Figure 9 shows that more communities are lost when switching to UNFL than to MULTI. While this may seem counterintuitive given that MULTI is more restrictive, this result underscores an important limitation of network flattening. Indeed, the process of flattening can lead to the loss of crucial network information and the introduction of noise in the network topology. This highlights the value of the MULTI approach, which preserves the structure and information inherent in individual layers, unlike UNFL. These findings also call into question the efficacy of the commonly used UNFL strategy, which is prevalent in the literature [31, 26, 52, 2, 30].
(c) UNFL vs. HST
(d) MULTI vs. HST
(e) UNFL vs. MEN
(f) MULTI vs. MEN
(g) UNFL vs. RPL
(h) MULTI vs. RPL
(i) UNFL vs. RTW
(j) MULTI vs. RTW
(k) UNFL vs. URL
(l) MULTI vs. URL
Figure 11: Lost, common, gained nodes characterization for UNFL and MULTI. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
Union flattening. Figure 9(a) provides insights on the results of UNFL with respect to INDI. Hence, UNFL exhibits similar patterns with blocks of common and gained communities and relatively few lost ones with respect to HST, MEN, and RTW. However, it shows a higher number of lost and gained communities compared with URL, suggesting that its information is poorly captured in the flattened network. Finally, UNFL has two communities in common with RPL and no lost communities but a significant number of gained ones, indicating that the detected communities differ substantially from those in the monomodal approach.
Multiplex community detection. Figure 9(b) demonstrates that MULTI has very few lost communities with respect to all monomodal approaches, with the exception of the URL. The transition from the co-reply layer to the multimodal approach reveals a higher number of gained communities, while other co-actions primarily show common communities and fewer gained ones. This suggests that communities identified using the multimodal MULTI approach retain much of the information from the monomodal approach while incorporating additional insights from the interplay between layers.
Discussion. This analysis highlights the effectiveness of MULTI in preserving existing information while uncovering new coordinated communities. Meanwhile, the UNFL is more prone to losing communities, further underscoring its limitations. The results suggest that MULTI is more effective in capturing the underlying structure and information of the network, providing a more comprehensive view of coordinated behavior. The node-level analysis reveals similar trends, with MULTI outperforming the UNFL in retaining and identifying nodes. This comparison underscores the importance of selecting an appropriate multimodal approach to effectively capture the complexity and richness of coordinated behavior in multiplex networks.
5.2.2 Multimodal communities characterization
To assess the effectiveness of multimodal approaches in identifying central nodes, we analyze the lost, common, and gained nodes defined in Section 3.3.3 and shown in Figure 10. This evaluation aims to determine whether multimodal methods retain or lose key influential nodes. Losing peripheral nodes is not critical, but missing highly connected or bridging nodes would indicate a limitation. Conversely, successfully preserving or discovering new influential nodes would highlight the advantage of multimodality in capturing a more complete network structure. To conduct this analysis, we compute key centrality and structural metrics for each node including degree centrality, eigenvector centrality, local clustering coefficient, and PageRank [75, 76]. The comparison is performed for both UNFL and MULTI. Figure 11 illustrates the distribution of these metrics across lost, common, and gained nodes for each co-action. We applied the Brunner-Munzel test [77] to compare distributions across the lost, common, and gained node subsets for each metric. Unlike the Mann-Whitney U test [78], the Brunner-Munzel test does not assume equal variances or distribution shapes, making it more robust to ties, outliers, and variance differences, and thus more suitable for our analysis. The comparison highlights the performance of both the UNFL and MULTI, offering insights into their ability to retain or identify influential nodes.
Union flattening. Lost and gained nodes are characterized by a higher degree centrality and eigenvector centrality switching from HST, MEN, RTW, and URL to UNFL, with respect to the common nodes. The unique exception is RPL, where the common nodes have lower values compared to gained and lost nodes. Instead, local clustering coefficient distribution is more homogeneous among the different co-actions. Finally, the page rank is higher for the lost and gained nodes, with respect to the common nodes, for HST, MEN, and RTW. Instead, for RPL and URL, the common nodes have higher values, but with similar distributions to the lost and gained ones.
Therefore, in case of HST, MEN, RTW, and URL, UNFL seems to lose the most important nodes, which are the most central and influential in the network, but at the same time it is able to detect new nodes that are not detected by the monomodal approach and that are as well central and influential. Instead, RPL shows a different pattern, where the common nodes are the most central and influential, while the lost and gained nodes are less central and influential, suggesting that both the monomodal approach and the UNFL are able to capture the most important nodes in the network.
Multiplex community detection. Observing MULTI, the common nodes have higher values of degree centrality and eigenvector centrality, with respect to the lost and gained nodes, for all the co-actions, with the exception of URL, where the lost nodes have higher values. Local clustering coefficient shows similar average for the lost, common, and gained nodes, for MEN, but with a higher variance for the lost and gained nodes. HST, RPL, and RTW show higher values for the common nodes, with respect to the lost and gained nodes. Finally, page rank is on average similar for all the co-actions and for the lost, common, and gained nodes, but with a higher variance for the common nodes.
Hence, degree and eigenvector centrality suggest that MULTI is able to detect the most central and influential nodes in the network, which are the same detected by the monomodal approach. Also local clustering coefficient suggests that on average the nodes detected by MULTI have similar distributions to the one lost.
Discussion. From the perspective of node centrality, the MULTI approach yields results that are generally consistent with those of the individual layers. Previous analyses showed that it retains most of the key nodes and communities from each modality, and although it tends to detect additional ones, these are typically less central. In contrast, UNFL exhibits the opposite behavior: it both gains and loses a notable number of nodes and communities, but crucially, it often discards highly central ones while still identifying new, influential actors. This suggests a significant misalignment between the results of UNFL and those of the individual layers. As such, these findings may raise concerns about the reliability of UNFL, especially in terms of preserving meaningful structural information.
6 Discussion and conclusions
We explored the role of multimodality in the detection and analysis of coordinated behavior, addressing its challenges and opportunities. To tackle this complexity, we introduced a novel methodology that integrates multiple types of user actions into a unified analytical framework for detecting multimodal coordinated behavior. By comparing monomodal approaches with flattened representations, and the multiplex community detection, our analysis highlighted the strengths and weaknesses of each approach, offering valuable insights into the detection and characterization of coordination in social media.
6.1 RQ1: Monomodal contributions
The first research question investigates how different monomodal approaches contribute to the analysis of coordinated behavior. Our findings suggest that there is no universally dominant modality. As shown in Section 5.1.2, RTW and MEN exhibit a high normalized mutual information score and identify fully overlapping communities, indicating strong agreement between these co-action layers. In contrast, the comparison between RTW and HST reveals a partial overlap—some communities are shared, while others are lost—highlighting moderate complementarity. Finally, RTW shares no communities with RPL and only one with URL, demonstrating substantial divergence in the coordinated structures each layer captures. Beyond metrics such as actor and edge coverage, which offer some anticipatory insights, it remains challenging to determine in advance which co-actions best reveal coordination. These findings underscore the importance of incorporating multiple co-action types to ensure that no significant form of coordination is overlooked. Additionally, Section 5.1.3 shows that even when monomodal approaches detect the same communities, these do not necessarily share similar structural characteristics. This is evident for RTW and URL, less so for RTW and HST, and absent for RTW and MEN. These results further emphasize the value of a multimodal framework, capable of integrating diverse and complementary signals of coordination.
6.2 RQ2: Multimodal vs. monomodal and different operationalizations
The second research question examines the contribution of multimodal approaches compared to monomodal ones in analyzing coordinated behavior. Appendix A shows that UNFL implemented with different flattening strategies produces similar outcomes, suggesting that the specific method of flattening has limited influence on the results. However, Section 5.2.1 demonstrates that the act of network flattening itself significantly impacts the detection of coordinated communities. In particular, the UNFL approach shares only a subset of communities with RTW, MEN, and HST, while failing to capture nearly all of those identified by URL. This indicates a substantial loss of information contributed by each monomodal layer. In contrast, the MULTI approach shows greater robustness: it retains almost all communities detected by monomodal methods and also uncovers additional coordinated structures. These findings raise concerns about the use of network flattening in the literature and advocate for the adoption of more expressive, though complex, multimodal methods. Section 5.2.2 further supports the advantage of MULTI, highlighting its ability to preserve influential nodes while also identifying new ones. By comparison, UNFL tends to lose highly central nodes, although it occasionally detects new, equally important ones. Overall, MULTI proves to be a more reliable approach for capturing both the breadth and depth of coordinated behavior.
6.3 Limitations and future works
This work presents the first comparative analysis of methods to opeartionalize the detection off multimodal coordinated behavior. Indeed, we define multiple operationalizations of multimodality, assessing the contributions of individual monomodal approaches, and evaluating the benefits of a multimodal framework over monomodal ones. The primary limitation of this study is that the analysis was conducted on a single dataset. As a result, the generalizability of our findings remains uncertain, and the conclusions drawn may be specific to the characteristics and dynamics of the observed case [79]. Future work will need to validate the proposed methodology across multiple datasets, ideally covering diverse social phenomena and contexts, to assess the robustness and consistency of our conclusions. Such efforts will be crucial to determine whether the observed patterns hold more broadly and to establish the general applicability of our multimodal framework. Given the lack of prior literature on multimodal coordination, we addressed several foundational research questions, necessarily leaving out additional operationalizations of multimodality. For instance, exploring alternative community detection algorithms, such as multilayer clique percolation [80] and ABACUS [81], could introduce new levels of multimodality within the intrinsic trade-off of coordination analysis (see Figure 2). Furthermore, future work could refine the integration across layers by systematically investigating the $\omega$ parameter of the Generalized Louvain algorithm, which controls the strength of inter-layer connections between the same user across modalities. Our current study focuses on five common co-actions, but the framework could be expanded to incorporate additional modalities, such as multimedia content sharing, which plays a central role on many social media platforms.
Finally, combining this approach with methods that account for temporal dynamics would yield deeper insights into how coordinated behavior evolves over time [33]. This would introduce the additional challenge of integrating temporal and multiplex network representations – an ambitious yet promising direction for future research.
In conclusion, we provide a solid foundation for understanding multimodal coordinated behavior, advancing both methodological and practical approaches to its detection and analysis. By offering a flexible and expandable framework, this work contributes to the broader effort of uncovering coordinated activities in social media and opens the way for future research in this domain.
Acknowledgements
This work was partly supported by SoBigData.it which receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021.; and by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; and by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006); and by the ERC project DEDUCE (Data-driven and User-centered Content Moderation) under grant #101113826; and by the HORIZON Europe projects TANGO - Grant Agreement n. 101120763; and by the European Commission under the NextGeneration EU programme – National Recovery and Resilience Plan (PNRR), under agreements: PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - "FAIR - Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI".
Appendix A Heatmap overlapping matrix of the communities
(a) RTW vs. INDI.
(b) MULTI vs. INDI.
Figure A.1: Heatmap of the overlapping matrix RTW and MULTI vs. INDI.
(a) UNFL (nw) vs. INDI.
(b) UNFL (ec) vs. INDI.
(c) UNFL (sum) vs. INDI.
Figure A.2: Heatmap of the overlapping matrix UNFL vs. INDI.
[…]